# Experimental demonstration of spontaneous symmetry breaking with emergent multi-qubit entanglement

### v2

@@ Abstract
This longitudinal correlation is equal to 0 for the $Z_{2}$-symmetric eigenstate $|X\rangle$, and equals to 1 for any combination of the $Z_{2}$-symmetry-breaking eigenstates $\left|g_{j}\right\rangle^{\otimes 6}$ and $\left|e_{j}\right\rangle^{\otimes 6}$, independent of their quantum coherence. This implies that ${\cal C}_{2}^{L}$ can be used as a measure to characterize the SSB degree. For the ideal LMG model, ${\cal C}_{2}^{L}$ would monotonously increase when $\Omega$ is adiabatically decreased, tending to 1 when $\Omega/\lambda\to 0$. Figure 3(c) displays ${\cal C}_{2}^{L}$ measured for different quench times. As expected, with the decrease of $\Omega$, it is progressively increased until $\tau=150$ ns. Due to the experimental imperfections, the maximum of ${\cal C}_{2}^{L}$ is limited to 0.67. The populations of $\left|g_{j}\right\rangle^{\otimes 6}$ and $\left|e_{j}\right\rangle^{\otimes 6}$ corresponding to this maximum are 0.34 and 0.32, respectively.
@@
where $U(\theta,\varphi)=\exp\left[-i\theta/2\left(\sin\varphi\sigma_{j}^{x}-\cos\varphi\sigma_{j}^{y}\right)\right]$ and $\Pi=\otimes_{j=1}^{6}(1-\sqrt{3}\sigma_{j}^{z})$. To measure the Wigner function at $(\theta,\varphi)$, the joint state readout is performed after a $\theta-$rotation around the axis with an angle $\varphi-\pi/2$ to the x-direction on the equatorial plane. The Wigner function reconstructed at $\tau=0$, 50, 100, and 150 ns are displayed in Fig. 4. The Wigner function corresponds to a quasi-probability distribution of the multi-qubit system on a unit sphere. The regions where the values are negative reflect the quantum interference effects, and the negativity serves as a signature of multi-qubit cat states [45]. The results clearly verify that the two symmetry-breaking state components are superimposed with each other during the SSB process, which represents one of the most remarkable features that distinguish phase transitions of finite-size quantum many-body systems from those of infinite-size ones. Such a signature is crucial to understand the quantum-to-classical transition in many-body systems, but has not been demonstrated so far.
@@
In conclusion, we have presented a demonstration of the SSB in a multi-qubit LMG model, engineered with a superconducting processor. The qubits are effectively coupled to each other through virtual photon exchange, mediated by a bus resonator dispersively coupled to the qubits. By tuning the competition between the external continuous driving and intra-qubit interaction, we can move the system from the $Z_{2}$-symmetric phase to the symmetry-breaking phase, where the two degenerate highest eigenstates of the governing Hamiltonian, superimposed with each other, form a GHZ-like cat state. We characterize the SSB extent with the average two-qubit longitudinal correlation. The quantum nature of the SSB phase is confirmed by the measured multi-qubit quantum correlation, as well as by the reconstructed multi-qubit Wigner function, which exhibits interference fringes with alternating positive and negative values. Besides fundamental interest, the demonstrated SSB process can be used for noise-insensitive many-body entanglement manipulation.
We thank Yi-Zhuang You at UC San Diego for the helpful discussion. This work was supported by the National Natural Science Foundation of China (Grants No. 12274080, No. 12204105, No. T2322030, No. T2121001, and No. 92065114), the Beiiing Nova Program (GrantNo.20220484121), and the Innovation Program for Quantum Science and Technology (Grants No. 2021ZD0300200 and No. 2021ZD0301800).

### v3

@@ Abstract
This longitudinal correlation is equal to 0 for the $Z_{2}$-symmetric eigenstate $|X\rangle$, and equals to 1 for any combination of the $Z_{2}$-symmetry-breaking eigenstates $\left|g_{j}\right\rangle^{\otimes 6}$ and $\left|e_{j}\right\rangle^{\otimes 6}$, independent of their quantum coherence. This implies that ${\cal C}_{2}^{L}$ can be used as a measure to characterize the SSB degree. For the ideal LMG model, ${\cal C}_{2}^{L}$ would monotonously increase when $\Omega$ is adiabatically decreased, tending to 1 when $\eta\to\infty$. Figure 3(c) displays ${\cal C}_{2}^{L}$ measured for different times. As expected, with the decrease of $\Omega$, it is progressively increased until $\tau=149$ ns. Due to the experimental imperfections, the maximum of ${\cal C}_{2}^{L}$ is limited to 0.82.
@@
where $U(\theta,\varphi)=\exp\left[-i\theta/2\left(\sin\varphi\sigma_{j}^{x}-\cos\varphi\sigma_{j}^{y}\right)\right]$ and $\Pi=\otimes_{j=1}^{6}(1-\sqrt{3}\sigma_{j}^{z})$. To measure the Wigner function at $(\theta,\varphi)$, the joint state readout is performed after a $\theta-$rotation around the axis with an angle $\varphi-\pi/2$ to the x-direction on the equatorial plane. The Wigner function reconstructed at $\tau=0$, 50, 100, and 149 ns are displayed in Fig. 4. The Wigner function corresponds to a quasi-probability distribution of the multi-qubit system on a unit sphere. The regions where the values are negative reflect the quantum interference effects, and the negativity serves as a signature of multi-qubit cat states [51]. These results clearly verify that the two symmetry-breaking state components are superimposed with each other during the QPT. We note that this quantum-mechanical feature vanishes when the qubit number tends to infinity due to the linear scaling of the decoherence rates (see Sec. S7 of the Supplemental Material [45] for details), thereby distinguishing phase transitions of finite-size quantum many-body systems from those of infinite-size ones.
@@
In conclusion, we have presented a demonstration of the SSB in the QPT of the multi-qubit LMG model, engineered with a superconducting processor. The qubits are effectively coupled to each other through virtual photon exchange, mediated by a bus resonator dispersively coupled to the qubits. By tuning the competition between the external continuous driving and intra-qubit interaction, we can move the system from the $Z_{2}$-symmetric phase to the symmetry-breaking phase, where the two degenerate highest eigenstates of the governing Hamiltonian, superimposed with each other, form a GHZ-like cat state. We characterize the SSB extent with the average two-qubit longitudinal correlation. The quantum nature of the SSB phase is confirmed by the measured multi-qubit quantum correlation, as well as by the reconstructed multi-qubit Wigner function, which exhibits interference fringes with alternating positive and negative values. Besides fundamental interest, the demonstrated SSB process can be used for noise-insensitive many-body entanglement manipulation.
Acknowledgements.
We thank Yi-Zhuang You at UC San Diego for the helpful discussion. This work was supported by the National Natural Science Foundation of China (Grants No. 12274080, No. 12474356, No. 12475015, No. 12204105, No. T2322030, No. T2121001, and No. 92065114), the Beiiing Nova Program (Grant No. 20220484121), and the Innovation Program for Quantum Science and Technology (Grants No. 2021ZD0300200 and No. 2021ZD0301800).